# Human-Robot Interaction Conversational User Enjoyment Scale (HRI CUES)

### v2

@@ VII Conclusion
Our research contributed a novel scale for measuring user enjoyment in conversations with a robot from an external perspective. The scale was developed through rigorous discussions of three annotators with complementary and relevant backgrounds to user enjoyment and the application domain. Older adults’ interactions with a companion robot were used as the basis for developing the scale, which was evaluated on 174 minutes of interactions of 25 participants. Inter-rater reliability analysis showed the importance of using multiple annotators, with moderate to good alignment, where the disagreements arose from the complexity and the subjectivity of user enjoyment, where a user can show various signs of enjoyment and dis-enjoyment within a single conversation exchange. The overall user enjoyment rated per interaction correlated significantly with users’ perceived level of strangeness of the conversation, which signifies that the (dis)comfort experienced in the interaction was correctly identified by the annotators, and shows the importance of including dis-enjoyment levels in the scale. These findings validate our user enjoyment scale and emphasize the critical role of methodological rigor in assessing subjective experiences within conversational robot interactions. Our study emphasizes the value of using multiple annotators and proposes potential scale refinements to further enhance consistency in quantifying the nuanced concept of enjoyment across application domains. The developed scale and the corresponding dataset aim to provide a tool for measuring user enjoyment from an external perspective to supplement self-reported user enjoyment responses in HRI research, with future potential application for autonomous detection of user enjoyment in real-time in robots and agents for adapting conversations contingently to provide enjoyable and long-lasting interactions.

### v3

@@ VII Conclusion
Our research contributed a novel scale for measuring user enjoyment in conversations with a robot from an external perspective. The scale was developed through rigorous discussions of three annotators with complementary and relevant backgrounds to user enjoyment and the application domain. Older adults’ interactions with a companion robot were used as the basis for developing the scale, which was evaluated on 174 minutes of interactions of 25 participants. Inter-rater reliability analysis showed the importance of using multiple annotators, with moderate to good alignment, where the disagreements arose from the complexity and the subjectivity of user enjoyment, especially when a user shows various signs of enjoyment and dis-enjoyment within a single conversation exchange. The overall user enjoyment rated per interaction correlated significantly with users’ perceived level of strangeness of the conversation, which signifies that the (dis)comfort experienced in the interaction was correctly identified by the annotators, and shows the importance of including dis-enjoyment levels in the scale. Using our scale with LLMs has been shown to correlate even more strongly with users’ self-reported ratings, paving a promising path toward automation. These findings support that our user enjoyment scale is a viable measure for external assessment of user enjoyment, and emphasize the critical role of methodological rigor in assessing subjective experiences within conversational robot interactions. Our study emphasizes the value of using multiple annotators and proposes potential scale refinements to further enhance consistency in quantifying the nuanced concept of enjoyment across application domains. The developed scale and the corresponding dataset aim to provide a tool for measuring user enjoyment from an external perspective to supplement self-reported user enjoyment responses in HRI research, with future potential application for autonomous detection of user enjoyment in real-time in robots and agents for adapting conversations contingently to provide enjoyable and long-lasting interactions.